How much electrical potential energy does a proton lose as it falls through a
potential drop of 5kV?
1
Expert's answer
2020-06-29T14:03:24-0400

Explanations & Calculations


  • Energy variation/work done of a charge ("q" ) moves through a voltage difference ("\small \Delta V" ) is given by

"\qquad\qquad
\begin{aligned}
\small W = q\Delta V
\end{aligned}"

  • This energy is either stored or released according to the moving direction.
  • Charge of a proton is the same to that of an electron but opposite in polarity.
  • q proton = "\small +1.6 \times 10^{-19} C"


  • Therefore, potential energy drop of the proton is,

"\qquad\qquad
\begin{aligned}
\small W&= \small 1.6 \times 10^{-19} C \times(5\times 10^3 V)\\
&= \small \bold{8\times10^{-16} J}
\end{aligned}"